OpenStudy (anonymous):

log(4x-2)-log(x-3)=1

OpenStudy (alfie):

\[\log \frac{4x-2}{x-3} = loge\] Or log10, if you are using base 10 logarithms. You can now do (4x-2)/(x-3) = e (or 10) and solve for x. Bye!

OpenStudy (anonymous):

thanks!